can you please help me do these problems?
solve by graphing method.
1)y - x = 4
y = 3x + 2

2)x + y = 1
5x + y =-7
3)4x + 2y =6
x - y =3

solve by the substitue method
4)
a=3b
a-5b=16


5)
8c-d=-3
4c+5d=15


6)
9p=2q-6
3p-q=12





Answer by Alan3354(21605)

can you please help me do these problems?
solve by graphing method.
1)y - x = 4
y = 3x + 2
2)x + y = 1
5x + y =-7
3)4x + 2y =6
x - y =3
For graphs, go to http://www.padowan.dk.com/graph/ and DL the free software.
solve by the substitue method
4)
a=3b
a-5b=16
----------
Sub for a into eqn 2
3b-5b = 16
-2b = 16
b = -8
a = -24
-------------------
5)
8c-d=-3 --> d = 8c+3
4c+5d=15
Sub for d into 2nd eqn
4c + 5(8c+3) = 15
4c + 40c+15 = 15
44c = 0
c = 0
d = 3
-----------
6)
9p=2q-6
3p-q=12
Do it like the 2 above.